To obtain the distribution constant of iodine between water and carbon tetrachloride (CCI,), two
solutions were placed together to attain the equilibrium: 50 mL of a saturated iodine solution in CCI,
+ 150 mL of water.

After 1 hour shaking at 25°C, the equilibrium was achieved. Two portions of each solution were
sampled and titrated with sodium thiosulphate. 15 mL of the organic phase needed 17,0 mL of
thiosulphate, and 50 mL of the aqueous solution needed 22 mL of a thiosulphate solution 5 times
diluted. Obtain the distribution coefficient.
1
Expert's answer
2020-06-29T07:51:24-0400

Solution.

"K = \frac{C(I2 \ in \ H2O)}{C(I2 \ in \ CCl4)}"

"C(I2 \ in \ org \ ph) = \frac{17*C}{15} = 1.133 C"

"C(I2 \ in \ water \ ph) = \frac{22*C*0.2}{50} = 0.088 C"

"C1 = \frac{1.133C*50}{15} = 3.78 C"

"C2 = \frac{0.088C*150}{50} = 0.264 C"

"K = \frac{0.264C}{3.78C} = 0.070"

Answer:

K = 0.070